A project's base case or most likely NPV is $50,000, and assume its probability of

occurrence is 60%. Assume the best case scenario NPV is 60% higher than the

base case and assume the worst scenario NPV is 30% lower than the base case. Both

the best case scenario and the worst case scenario have a 20% probability of

occurrence. Find the project's coefficient of variation.
Answer #1

Expected NPV=Sum(probability*NPV)=50000*60%+50000*1.6*20%+50000*0.7*20%=53000.00

Standard deviation of NPV=sqrt(Sum(probability*(NPV-Expected NPV)^2))=sqrt(60%*(50000-53000)^2+20%*(50000*1.6-53000)^2+20%*(50000*0.7-53000)^2)=14696.93846

Coefficient of variation=Standard deviation/Expected NPV=14696.93846/53000=0.27730